OpenStudy (anonymous):

3. Simplify (4x + 1) + (–3x + 9) (1 point) x + 10 7x – 8 7x + 10 x – 8

OpenStudy (anonymous):

can you plz help me out with some problems?

OpenStudy (eyust707):

since there is a plus sign in between the parenthesis it doesnt really change much

OpenStudy (eyust707):



OpenStudy (anonymous):

I actually answered that one already can u help me with a few more PLZ

OpenStudy (eyust707):

sure

OpenStudy (anonymous):

Thank u sooo much

OpenStudy (anonymous):

Complete the table and choose the solution for each given value of x in questions 8–13. y = –x2 + 2

OpenStudy (anonymous):

8. When x = –2 (1 point) y = –(–2)2 + 2 = –2; (–2, –2) y = (–2)2 + 2 = 2; (–2, 2) y = –(2)2 + 2 = –2; (2, –2) y = –(–2)2 – 2 = 2; (–2, –2) 9. When x = – 1 (1 point) y = –(–1)2 + 1 = 2; (–1, 1) y = (–1)2 + 2 = 1; (–2, 1) y = –(–1)2 + 2 = 1; (–1, 1) y = (–1)2 + 2 = 1; (1, –1) 10. When x = 0 (1 point) y = –(0)2 + 2 = 2; (0, 2) y = –(0)2 + 2 = –2; (0, –2) y = (0)2 – 2 = 2; (2, 0) 11. When x = 1 (1 point) y = –(1)2 – 2 = 1; (1, 1) y = –(1)2 + 2 = 1; (1, 1) y = (1)2 + 2 = –1; (1, 1) y = (1)2 + 2 = 1; (1, –1) 12. When x = 2 (1 point) y = (2)2 – 2 = 2; (2, –2) y = –(2)2 – 2 = –2; (2, 2) y = (2)2 + 2 = –2; (2, 2) y = –(2)2 + 2 = –2; (2, –2)

OpenStudy (anonymous):

can u help me with these

OpenStudy (eyust707):

so for these all you do is plug the x value into the equation

OpenStudy (eyust707):

the equation then outputs a y value

OpenStudy (eyust707):

thats what goes in the y box

OpenStudy (anonymous):

can u give me an answer is u dont mind

OpenStudy (eyust707):

sry but thats actually against the rules

OpenStudy (eyust707):

i can teach you how to do it if you would like.. its actually pretty easy once you get the hang of it

OpenStudy (anonymous):

what rules

OpenStudy (eyust707):

Code of Conduct on the bottom left corner of your screen

OpenStudy (anonymous):

but the last time some one helped me they explained then gave me the answers

OpenStudy (eyust707):

not everyone follows the code of conduct

OpenStudy (eyust707):

I do

OpenStudy (eyust707):

here check it out its actually pretty easy

OpenStudy (anonymous):

well can u explain and then help me find the answer

OpenStudy (eyust707):

you have